OpenStudy (anonymous):

Solve using the square root method: 5x^2=60

OpenStudy (anonymous):

2√3 or -2√3

OpenStudy (rulnick):

x^2 = 60/5 = 12 ... then apply the square root method; take sqrt of each side: x = plus or minus sqrt(12)

OpenStudy (rulnick):

... = plus or minus 2 sqrt(3).

OpenStudy (anonymous):

my options are: A. 3.5,-3.5 B. 7.7,-7.7 C. 3.5 D. No real solution

OpenStudy (anonymous):

C. 3.5 (happy?) √12 = 3.46

OpenStudy (rulnick):

Problem is, a couple of people gave medals for the wrong answer. The answer is (A), 3.5, -3.5. See my solution above (and others).

OpenStudy (anonymous):

Thank you rulnick

OpenStudy (rulnick):

welcome